# Distractor-free Generalizable 3D Gaussian Splatting

### v2

@@ Abstract
We present DGGS, a novel framework that addresses the previously unexplored challenge: Distractor-free Generalizable 3D Gaussian Splatting (3DGS). It mitigates 3D inconsistency and training instability caused by distractor data in the cross-scenes generalizable train setting while enabling feedforward inference for 3DGS and distractor masks from references in the unseen scenes. To achieve these objectives, DGGS proposes a scene-agnostic reference-based mask prediction and refinement module during the training phase, effectively eliminating the impact of distractor on training stability. Moreover, we combat distractor-induced artifacts and holes at inference time through a novel two-stage inference framework for references scoring and re-selection, complemented by a distractor pruning mechanism that further removes residual distractor 3DGS-primitive influences. Extensive feedforward experiments on the real and our synthetic data show DGGS’s reconstruction capability when dealing with novel distractor scenes. Moreover, our generalizable mask prediction even achieves an accuracy superior to existing scene-specific training methods. Codes and models will be made publicly available.
@@ 1 Introduction
The widespread availability of mobile devices presents unprecedented opportunities for 3D reconstruction, fostering demand for feedforward 3D synthesis capabilities from casually captured images or video sequences (referred to as references). Recent approaches introduce generalizable 3D representations to address this challenge, eliminating per-scene optimization requirements, with 3D Gaussian Splatting (3DGS) demonstrating particular promise due to its efficiency charatan2024pixelsplat ; liu2025mvsgaussian ; chen2024mvsplat ; zhang2024transplat . In pursuit of scene-agnostic inference from references to 3DGS, existing methods simulate the complete pipeline ‘inputs references infer 3DGS and outputs novel query views’ within each training step, utilizing selected reference-query pairs while optimizing model parameters by query rendering losses.
Following this paradigm, a generalizable 3DGS requires both comprehensive training pairs and a trained scene-independent mechanism to understand 3D geometric correlations between references. However, these essential components face fundamental challenges from the distractor in unconstrained capture scenarios: (1) real-world scenes typically lack distractor-free data for training, and (2) distractor disrupt 3D consistency among limited references during training and inference processes.
Figure 1: Overview of Our Task. Distractors are unwanted transient objects in static scene reconstruction, such as buses, balloons, or anything. DGGS enables feedforward 3DGS reconstruction from limited distractor data while inferring corresponding distractor masks without extra supervision.
To address these problems, a straightforward solution is to integrate scene-specific distractor-free masking methods chen2024nerf ; ungermann2024robust ; sabour2024spotlesssplats into training loss of generalizable 3DGS, i.e. ‘MVSplat+’ in Sec. 5.2, enabling the prediction of distractor masks from the outliers of residual loss and actively suppressing them in loss function that would otherwise compromise training robustness sabour2023robustnerf . However, two main limitations emerge in this approach: (1) The existing loss-based masking strategies rely heavily on repeated optimization with sufficient inputs in a single scene and scene-specific hyperparameters sabour2023robustnerf . They are ineffective in our cross-scene generalizable training setting, where loss uncertainty increases due to inter-iteration scene transitions and uncertain scene hyperparameters. This uncertainty undermines the core assumption that high-loss regions correspond to distractor, potentially misclassifying the part of complex target areas as distractor masks in Fig.3 6 and resulting in inadequate supervision. (2) During feedforward inference, even when accurate masks are obtained, reconstruction in distractor areas continues to be affected by artifacts and remains incomplete due to limited references in Fig.7 10.
@@
For the first challenge, we design a Distractor-free Generalizable Training paradigm, incorporating a Reference-based Mask Prediction and a Mask Refinement module to enhance training stability based on precise generalizable distractor masking. Specifically, inspired by the observation in ‘references to 3DGS’ training paradigm that re-rendered reference non-distractor areas from 3DGS (inferred by references) are generally accurate and robust, we use these reference areas as guidance to filter out misclassified distractor regions in query view. This process is realized on the basis of the static geometric multi-view consistency of non-distractor areas. For higher mask accuracy, after decoupling the filtered masks into distractor and disparity error areas, our Mask Refinement module incorporates pre-trained segmentation results to fill distractor regions and a reference-based auxiliary loss to additionally supervise the unnoticed occluded regions in query view.
For the second challenge, which causes holes and artifacts in rendered views, we propose a two-stage Distractor-free Generalizable Inference framework. Specifically, in the first stage, we design a Reference Scoring mechanism based on the predicted coarse 3DGS and distractor masks from initially sampled references. These predicted scores guide the selection of minimal distractor and disparity references for fine reconstruction in the second stage. To further mitigate ghosting artifacts from residual distractor in the second stage, we introduce a Distractor Pruning strategy that eliminates distractor-associated gaussian primitives in the 3D space.
@@
Overall, we address a new task of Distractor-free Generalizable 3DGS as in Fig. 1, and this is, to our knowledge, the first work to explore this problem. For this objective, we present DGGS, a framework designed to alleviate distractor-related adverse effects during the training and inference phases of generalizable 3DGS. Extensive feedforward experiments on real distractor data have shown that our approach successfully enhances training robustness and improves artifacts and holes during inference while expanding cross-scene (outdoor scenes training and indoor scenes inference) generalizability in conventional scene-specific distractor-free models chen2024nerf ; sabour2023robustnerf ; Ren2024NeRF ; sabour2024spotlesssplats . Beyond real data, we also construct some synthetic distractor scenes based on Re10K and ACID datasets for further verification. Furthermore, our reference-based training paradigm achieves better generalizable distractor masking without any mask supervision, even outperforming scene-specific training distractor-free workschen2024nerf .
@@ 2.1 Generalizable 3D Reconstruction
Contemporary advances in generalizable 3D reconstruction seek to establish scene-agnostic representations, building upon early explorations in Neural Radiance Fields mildenhall2021nerf ; wang2021ibrnet ; wang2022attention ; liu2022neural ; bao2023insertnerf . However, these methods face significant bottlenecks due to their lack of explicit representations and rendering inefficiencies. The advent of 3D Gaussian Splatting kerbl20233d , an explicit representation optimized for efficient rendering, has sparked renewed interest in the field. Existing works involve inferring Gaussian primitive attributes from references directly and rendering in novel views. Analogous to NeRF-based approaches, 3DGS-related methods emphasize spatial comprehension from references, particularly focusing on depth estimation charatan2024pixelsplat ; chen2024mvsplat ; liu2025mvsgaussian ; zhang2024transplat ; liang2023gaufre . Subsequently, ReconX liu2024reconx and G3R chen2025g3r enhance reconstruction quality through the integration of additional video diffusion models and supplementary sensor inputs. The inherent reliance on high-quality references, however, makes generalizable reconstruction particularly susceptible to distractor, a persistent challenge in real-world applications. In this study, we discuss Distractor-free Generalizable 3DGS, an unexplored topic in the current works.
@@ 2.2 Scene-specific Distractor-free Reconstruction
It focuses on accurately reconstructing a specific static scene while mitigating the impact of distractor Ren2024NeRF (or transient objects sabour2023robustnerf ). As a pioneering work, NeRF-W martin2021nerf introduces additional embeddings to represent and eliminate transient objects in unstructured photo collections. Following a similar setting, subsequent works focus on mitigating the impact of distractor at the image level, which can generally be categorized into knowledge-based and heuristics-based methods.
Knowledge-based methods predict distractor using external knowledge sources. Among them, pre-trained features from ResNet zhang2024gaussian ; xu2024wild , diffusion models sabour2024spotlesssplats , and DINO Ren2024NeRF ; kulhanek2024wildgaussians guide visibility map prediction, effectively weighting reconstruction loss. Recent works chen2024nerf ; otonari2024entity ; nguyen2024rodus directly employ state-of-the-art segmentation models such as SAM kirillov2023segment or Entity Segmentation qi2022high to establish clear distractor boundaries. Although these approaches demonstrate certain improvements martin2021nerf ; chen2022hallucinated ; lee2023semantic with additional priors, they struggle to differentiate the distractor from static target scenes chen2024nerf ; otonari2024entity .
@@
Heuristics-based approaches employ handcrafted statistical metrics to distinguish distractor, emphasizing robustness and uncertainty analysis sabour2023robustnerf ; goli2024bayes ; ungermann2024robust . These methods exploit the observation that regions containing distractor typically manifest optimization inconsistencies. Therefore, they seek to predict outliers and mitigate their impact in residual losses. Regrettably, these approaches suffer from significant scene-specific data dependencies and frequently confound distractor with inherently challenging reconstruction regions, limiting their effectiveness in generalizable contexts.
@@
Recently, there has been increasing advocacy for integrating the two methods mentioned above otonari2024entity ; chen2024nerf . Entity-NeRF otonari2024entity integrates an existing Entity Segmentation qi2022high and extra entity classifier to determine distractor among entities by analyzing the rank of residual loss. Similarly, NeRF-HuGS chen2024nerf integrates pre-defined Colmap and Nerfacto tancik2023nerfstudio for capturing high and low-frequency features of static targets, while using SAM kirillov2023segment to predict clear distractor masks. However, in our settings, acquiring additional entity classifiers or employing pre-defined scene-level knowledge such as Colmap and Nerfacto is nearly impossible, and residual loss becomes unreliable compared to single-scene optimization due to the absence of iteratively refined representation. Moreover, with limited references in unseen scenes, despite obtaining distractor masks, traditional Distractor-free methods struggle to handle occluded regions and artifacts. Therefore, we present a novel Distractor-free Generalizable framework that jointly addresses distractor effects in training and inference phases.
@@ 3 Preliminaries
3DGS $\mathcal{G}$ represents 3D scenes by splatting numerous anisotropic gaussian primitives. Each gaussian primitive is characterized by a set of attributes $\mathbb{A}$, including position $\bm{p}$, opacity $\alpha$, covariance matrix $\mathbf{\Sigma}$, and spherical harmonic coefficients for color $\hat{\bm{c}}$. To ensure positive semi-definiteness, the covariance matrix $\mathbf{\Sigma}$ is decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$, such that $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$. Consequently, the color value after splatting on view $\mathbf{P}$ is:
@@ 3.2 Generalizable 3DGS
where $\left\{\mathbf{P}_{i}\right\}^{N}_{i=1}$ and $\mathbf{P}_{T}$ are the camera extrinsics of the reference and query images, and $N$ denotes the number of references. The $\mathcal{F}$ denotes the process of projecting encoded 2D features into 3D space and refining chen2024mvsplat ; charatan2024pixelsplat . The 3D features are then decoded into the corresponding 3DGS attributes and rendered into query views, through which the network can be optimized by the query MSE loss.
@@ 3.3 Robust Masks for 3D Reconstruction
Unlike conventional controlled static environments, our in-the-wild scenarios contain not only static elements but also distractor sabour2023robustnerf , making it difficult to maintain 3D geometric consistency and training stability. Building on prior Scene-specific Distractor-free reconstruction research sabour2023robustnerf ; sabour2024spotlesssplats ; Ren2024NeRF ; chen2024nerf , we integrate the mask-based robust optimization loss in our pipeline (i.e. ‘MVSplat+ *’ in experiment, where * represents different mask prediction methods) that can predict and filter out distractor (outliers) in training process. Hence, Eq. 2 is modified, where $\odot$ denotes pixel-wise multiplication:
@@
Here, $\mathcal{M}_{Rob}$ represents the outlier masks on query view, where distractor is set to zero, while target static regions are set to one. Our work introduces a simple heuristic method sabour2023robustnerf as the foundation. The hyperparameters $\rho_{1}$ and $\rho_{2}$ are fixed across all scenes.
@@
where $\mathcal{C}$ represents the kernel operator and the $\rho_{1}$ as well as $\rho_{2}$ remain consistent with sabour2023robustnerf . Despite various mask refinements in follow-up studies otonari2024entity ; chen2024nerf ; sabour2024spotlesssplats , their heavy dependence on loss $\left\|\mathbf{I}_{T}-\mathcal{G}\left(\mathbf{P}_{T}\right)\right\|_{2}$ leads to extensive misclassification of difficult-to-feedforward-inference parts as distractor regions, as show in Fig. 3 and Fig. 6, which is addressed in subsequent sections.
@@ 4.1.1 Reference-based Mask Prediction
This Mask Prediction module aims to enhance the $\mathcal{M}_{Rob}$ accuracy within each iterative various scene, ensuring that optimization efforts remain focused on more non-distractor areas. It is essential for generalization training, since Eq. 4 inevitably misclassifies certain target regions as distractor, particularly those presenting challenges for feedforward inference, as in Fig. 3 6, which impedes the model’s comprehension of geometric 3D consistency in Fig. 5. Our inspiration stems from an intuitive observation: the 3DGS inferred from references maintains stable and accurate re-rendering results in non-distractor regions under reference views. Specifically, we introduce a mask filter that harnesses non-distractor regions from re-rendered references (i.e. from references to infer 3DGS and render back to reference views) to identify and remove falsely labeled distractor regions in $\mathcal{M}_{Rob}$ under query view, based on the multi-view consistency of non-distractor static objects. Given $i\in N$ denotes references and $\rho_{Ref}$ is a hyperparameter (as 0.001), discussed in Fig. 10, re-rendered reference non-distractor masks $\mathcal{M}_{Ref_{i}}$ and corresponding projected query masks $\mathcal{M}_{Qry_{i}}$ are
@@
However, given the inherent noise presence in $\mathcal{M}_{Ref_{i}}$, $\mathcal{M}_{Qry_{i}}$ exhibits limited precision. To solve this problem, we incorporate a pre-trained segmentation model for mask filling, while designing a multi-mask fusion strategy to counteract noise-induced deviations. Following chen2024nerf ; otonari2024entity , we incorporate a state-of-the-art Entity Segmentation Model qi2022high to refine $\mathcal{M}_{Ref_{i}}$ into $\mathcal{M}^{En}_{Ref_{i}}$, which has the capability to fill different entities when the predicted distractor-region exceeds the threshold of total pixel count for that entity. After substituting $\mathcal{M}_{Ref_{i}}$ with $\mathcal{M}^{En}_{Ref_{i}}$ in Eq. 5, we use an intersection operation to fuse all (N) $\mathcal{M}_{Qry_{i}}$, then filter $\mathcal{M}_{Rob}$ based on it, obtaining the reference-based mask $\mathcal{M}_{Q}$, which can filter out misclassified regions in $\mathcal{M}_{Rob}$ to some extent, as shown in Fig. 3.
$\mathcal{M}_{Q}=\left\{\bigcap_{i=1}^{N}\mathcal{M}_{Qry_{i}}\right\}\bigcup\mathcal{M}_{Rob}.$ (6)
@@
Figure 3: The Mask Evolution in Sec. 4.1. $\mathcal{M}_{Q}$ is obtained by filtering $\mathcal{M}_{Rob}$ from the references non-distractor regions, which is then filled by decoupling $\mathcal{M}_{D}$ and using segmentation results to get final $\mathcal{M}$ as Eq. 4 6 7 8. Without references filter, target regions are often misidentified as distractor.

### v3

@@ Abstract
We present DGGS, a novel framework that addresses the previously unexplored challenge: Distractor-free Generalizable 3D Gaussian Splatting (3DGS). Previous generalizable 3DGS works are often limited to static scenes, struggling to mitigate distractor impacts in training and inference phases, which leads to training instability and inference artifacts. To address this new challenge, we propose a distractor-free generalizable training paradigm and corresponding inference framework, which can be directly integrated into existing Generalizable 3DGS frameworks. Specifically, in our training paradigm, DGGS proposes a feed-forward mask prediction and refinement module based on the 3D consistency of references and semantic prior, effectively eliminating the impact of distractor on training loss. Based on these masks, we combat distractor-induced artifacts and holes at inference time through a novel two-stage inference framework for reference scoring and re-selection, complemented by a distractor pruning mechanism that further removes residual distractor 3DGS-primitive influences. Extensive feed-forward experiments on the real and our synthetic data show DGGS’s reconstruction capability when dealing with novel distractor scenes. Moreover, our feed-forward mask prediction even achieves an accuracy superior to scene-specific Distractor-free methods.
@@ 1 Introduction
The widespread availability of mobile devices presents unprecedented opportunities for 3D reconstruction, fostering demand for feed-forward 3D synthesis capabilities from casually captured images or video sequences (referred to as references). Recent approaches introduce generalizable 3D representations to address this challenge, eliminating per-scene optimization requirements, with 3D Gaussian Splatting (3DGS) demonstrating particular promise due to its efficiency (Charatan et al., 2024; Liu et al., 2025; Chen et al., 2024b; Zhang et al., 2024a). In pursuit of scene-agnostic inference from references to 3DGS, existing methods project reference features onto 3D space to predict 3DGS attributes and simulate the complete pipeline: input references, infer 3DGS, and render novel query views, within each training step. This process utilizes selected reference-query pairs for training and optimizes the model to learn the reference-query 3D consistency through query rendering losses.
While promising, this paradigm faces two major challenges in real-world, unconstrained capture scenarios due to the presence of distractor (e.g., transient objects such as vehicles or pedestrians). First, during training, real-world data often contains distractor that disrupt 3D consistency, limiting training to confined static scenes. Second, during inference, distractor in the reference images cannot be properly projected into 3D space, resulting in unwanted artifacts in the reconstructed 3D scene.
To tackle these issues, we propose Distractor-free Generalizable 3D Gaussian Splatting (DGGS), a novel framework that enhances training stability when training generalizable 3DGS models under distractor-data and mitigates distractor-induced artifacts during the inference process. This framework builds on two key components: a Distractor-free Generalizable Training paradigm and a Distractor-free Generalizable Inference framework. The core idea behind them lies in the discussion about how to predict distractor masks in a feed-forward manner and use them for training and inference. Unlike existing scene-specific distractor-free masking methods that require sufficient input and iterative optimization (Chen et al., 2024a; Ungermann et al., 2024; Sabour et al., 2024), our method takes advantage of the inherent 3D consistency across references to infer distractor masks in each training iteration. These masks are then applied to exclude distractor regions from reconstruction loss during training, improving stability, and are further used during inference to prioritize cleaner references and suppress artifacts.
@@
Specifically, in Distractor-free Generalizable Training paradigm, we design a Reference-based Mask Prediction and a Mask Refinement module for generalizable distractor masking, which is based on a core observation, re-rendered reference non-distractor areas from 3DGS (inferred by references) are generally accurate and robust. Therefore, given an initialized mask, we use these non-distractor areas in reference as guidance to filter out misclassified distractor regions in query. This process relies on the static geometric multi-view consistency of non-distractor areas. For higher mask accuracy, after decoupling the filtered masks into distractor and disparity error areas, our mask refinement module incorporates pre-trained segmentation results to fill distractor regions and designs a reference-based auxiliary loss to additionally supervise the unnoticed occluded regions in query view. Based on these masks, we propose a two-stage Distractor-free Generalizable Inference framework for mitigating holes and artifacts at inference time. In the first stage, we introduce more candidate references and design a Reference Scoring mechanism to score candidator based on the predicted distractor masks. These predicted scores guide the references selection with minimal distractors and disparity for fine reconstruction in the second stage. To further mitigate ghosting artifacts from residual distractor in the second stage, we introduce a Distractor Pruning strategy that eliminates distractor-associated 3D gaussian primitives.
Figure 1: Overview of Our Task. Distractors are unwanted transient objects in static scene reconstruction, such as buses, balloons, or anything. DGGS enables feed-forward 3DGS reconstruction from limited distractor data while inferring corresponding distractor masks without extra supervision.
@@
Overall, we address a new task of Distractor-free Generalizable 3DGS as in Fig. 1, and this is, to our knowledge, the first work to explore this problem. For this objective, we present DGGS, a framework designed to alleviate distractor-related adverse effects during the training and inference phases of generalizable 3DGS. Extensive feed-forward experiments on real distractor data have shown that our approach successfully enhances training robustness and improves artifacts and holes during inference while expanding cross-scene (outdoor scenes training and indoor scenes inference) generalizability in conventional scene-specific distractor-free models (Chen et al., 2024a; Sabour et al., 2023; Ren et al., 2024; Sabour et al., 2024). Beyond real data, we also construct some synthetic distractor scenes based on Re10K and ACID datasets for further verification. Furthermore, our reference-based training paradigm achieves better generalizable distractor masking without any mask supervision, even outperforming scene-specific training distractor-free works(Chen et al., 2024a).
@@ 2.1 Generalizable 3D Reconstruction
Contemporary advances in generalizable 3D reconstruction seek to establish scene-agnostic representations, building upon early explorations in Neural Radiance Fields (Mildenhall et al., 2021; Wang et al., 2021; 2022; Liu et al., 2022; Bao et al., 2023a). However, these methods face significant bottlenecks due to their lack of explicit representations and rendering inefficiencies. The advent of 3D Gaussian Splatting (Kerbl et al., 2023), an explicit representation optimized for efficient rendering, has sparked renewed interest in the field. Existing works involve inferring Gaussian primitive attributes from references directly and rendering in novel views. Analogous to NeRF-based approaches, 3DGS-related methods emphasize spatial comprehension from references, particularly focusing on depth estimation (Charatan et al., 2024; Chen et al., 2024b; Liu et al., 2025; Zhang et al., 2024a; Liang et al., 2023). Subsequently, ReconX (Liu et al., 2024) and G3R (Chen et al., 2025) enhance reconstruction quality through the integration of additional video diffusion models and supplementary sensor inputs. The inherent reliance on high-quality references, however, makes generalizable reconstruction particularly susceptible to distractor, a persistent challenge in real-world applications. In this study, we discuss Distractor-free Generalizable 3DGS, an unexplored topic.
@@ 2.2 Scene-specific Distractor-free Reconstruction
It focuses on accurately reconstructing a specific static scene while mitigating the impact of distractor (Ren et al., 2024) (or transient objects (Sabour et al., 2023)). As a pioneering work, NeRF-W (Martin-Brualla et al., 2021) introduces additional embeddings to represent and eliminate transient objects in unstructured photo collections. Following a similar setting, subsequent works focus on mitigating the impact of distractor at the image level, which can generally be categorized into knowledge-based and heuristics-based methods.
@@
Knowledge-based methods predict distractor using external knowledge sources. Among them, pre-trained features from ResNet (Zhang et al., 2024b; Xu et al., 2024), diffusion models (Sabour et al., 2024), and DINO (Ren et al., 2024; Kulhanek et al., 2024) guide visibility map prediction, effectively weighting reconstruction loss. Recent works (Chen et al., 2024a; Otonari et al., 2024; Nguyen et al., 2024) directly employ state-of-the-art segmentation models such as SAM (Kirillov et al., 2023) or Entity Segmentation (Qi et al., 2022) to establish clear distractor boundaries. Although these approaches demonstrate certain improvements (Martin-Brualla et al., 2021; Chen et al., 2022; Lee et al., 2023) with additional priors, they struggle to differentiate the distractor from static target scenes (Chen et al., 2024a; Otonari et al., 2024). Heuristics-based approaches employ handcrafted statistical metrics to distinguish distractor, emphasizing robustness and uncertainty analysis (Sabour et al., 2023; Goli et al., 2024; Ungermann et al., 2024). These methods exploit the observation that regions containing distractor typically manifest optimization inconsistencies. Therefore, they seek to predict outliers and mitigate their impact in residual losses. Regrettably, these approaches suffer from significant scene-specific data dependencies and frequently confound distractor with inherently challenging reconstruction regions, limiting their effectiveness in generalizable contexts.
@@
Recently, there has been increasing advocacy for integrating the two methods (Otonari et al., 2024; Chen et al., 2024a). Entity-NeRF (Otonari et al., 2024) integrates an existing Entity Segmentation (Qi et al., 2022) and extra entity classifier to determine distractor among entities by analyzing the rank of residual loss. Similarly, NeRF-HuGS (Chen et al., 2024a) integrates pre-defined Colmap and Nerfacto (Tancik et al., 2023) for capturing high and low-frequency features of static targets, while using SAM (Kirillov et al., 2023) to predict clear distractor masks. However, in our settings, acquiring additional entity classifiers or employing pre-defined scene-level knowledge such as Colmap and Nerfacto is nearly impossible, and residual loss becomes unreliable compared to single-scene optimization due to the absence of iteratively refined representation. Moreover, with limited references in unseen scenes, despite obtaining distractor masks, traditional Distractor-free methods struggle to handle occluded regions and artifacts. Therefore, we present a novel Distractor-free Generalizable framework that jointly addresses distractor effects in training and inference phases.
@@ 3 Preliminaries
3DGS (Kerbl et al., 2023; Bao et al., 2025) $\mathcal{G}$ represents 3D scenes by splatting numerous anisotropic gaussian primitives. Each gaussian primitive is characterized by a set of attributes $\mathbb{A}$, including position $\bm{p}$, opacity $\alpha$, covariance matrix $\mathbf{\Sigma}$, and spherical harmonic coefficients for color $\hat{\bm{c}}$. To ensure positive semi-definiteness, the covariance matrix $\mathbf{\Sigma}$ is decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$, such that $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$. Consequently, the color value after splatting on view $\mathbf{P}$ is:
@@ 3.2 Generalizable 3DGS
where $\left\{\mathbf{P}_{i}\right\}^{N}_{i=1}$ and $\mathbf{P}_{T}$ are the camera extrinsics of the reference and query images, and $N$ denotes the number of references. The $\mathcal{F}$ denotes the process of projecting encoded 2D features into 3D space and refining (Chen et al., 2024b; Charatan et al., 2024). The 3D features are then decoded into the corresponding 3DGS attributes and rendered into query views, through which the network can be optimized by the query MSE loss.
@@ 3.3 Robust Masks for 3D Reconstruction
Unlike conventional controlled static environments, our in-the-wild scenarios contain not only static elements but also distractor (Sabour et al., 2023), making it difficult to maintain 3D geometric consistency and training stability. Building on prior Scene-specific Distractor-free reconstruction research (Sabour et al., 2023; 2024; Ren et al., 2024; Chen et al., 2024a), we integrate the mask-based robust optimization loss in our pipeline (i.e. ‘MVSplat+ *’ in experiment, where * represents different mask prediction methods) that can predict and filter out distractor (outliers) in training process. Hence, Eq. 2 is modified, where $\odot$ denotes pixel-wise multiplication:
@@
Here, $\mathcal{M}_{Rob}$ represents the outlier masks on query view, where distractor is set to zero, while target static regions are set to one. Our work introduces a simple heuristic method (Sabour et al., 2023) as the foundation. The hyperparameters $\rho_{1}$ and $\rho_{2}$ are fixed across all scenes.
@@
where $\mathcal{C}$ represents the kernel operator and the $\rho_{1}$ as well as $\rho_{2}$ remain consistent with (Sabour et al., 2023). Despite various mask refinements in follow-up studies (Otonari et al., 2024; Chen et al., 2024a; Sabour et al., 2024), their heavy dependence on loss $\left\|\mathbf{I}_{T}-\mathcal{G}\left(\mathbf{P}_{T}\right)\right\|_{2}$ leads to extensive misclassification of difficult-to-feed-forward-inference parts as distractor regions, as show in Fig. 3 and Fig. 6, which is addressed in subsequent sections.
@@ 4.1.1 Reference-based Mask Prediction
This Mask Prediction module aims to enhance the $\mathcal{M}_{Rob}$ accuracy within each iterative various scene, ensuring that optimization efforts remain focused on more non-distractor areas. It is essential for generalization training, since Eq. 4 inevitably misclassifies certain target regions as distractor, particularly those presenting challenges for feed-forward inference, as in Fig. 3 6, which impedes the model’s comprehension of geometric 3D consistency in Fig. 5. Our inspiration stems from an intuitive observation: the 3DGS inferred from references maintains stable and accurate re-rendering results in non-distractor regions under reference views. Specifically, we introduce a mask filter that harnesses non-distractor regions from re-rendered references (i.e. from references to infer 3DGS and render back to reference views) to identify and remove falsely labeled distractor regions in $\mathcal{M}_{Rob}$ under query view, based on the multi-view consistency of non-distractor static objects. Given $i\in N$ where $i$ denotes references and $\rho_{Ref}$ is a hyperparameter (set to 0.001), discussed in Fig. 10, re-rendered reference non-distractor masks $\mathcal{M}_{Ref_{i}}$ and corresponding projected query masks $\mathcal{M}_{Qry_{i}}$ are
@@
However, given the inherent noise presence in $\mathcal{M}_{Ref_{i}}$, $\mathcal{M}_{Qry_{i}}$ exhibits limited precision. To solve this problem, we incorporate a pre-trained segmentation model for mask filling, while designing a multi-mask fusion strategy to counteract noise-induced deviations. Following (Chen et al., 2024a; Otonari et al., 2024), we incorporate a state-of-the-art Entity Segmentation Model (Qi et al., 2022) to refine $\mathcal{M}_{Ref_{i}}$ into $\mathcal{M}^{En}_{Ref_{i}}$, which has the capability to fill different entities when the predicted distractor-region exceeds the threshold of total pixel count for that entity. After substituting $\mathcal{M}_{Ref_{i}}$ with $\mathcal{M}^{En}_{Ref_{i}}$ in Eq. 5, we use an intersection operation to fuse all (N) $\mathcal{M}_{Qry_{i}}$, then filter $\mathcal{M}_{Rob}$ based on it, $\mathcal{M}_{Q}=\left\{\bigcap_{i=1}^{N}\mathcal{M}_{Qry_{i}}\right\}\bigcup\mathcal{M}_{Rob}$, obtaining the reference-based mask $\mathcal{M}_{Q}$, which can filter out misclassified regions in $\mathcal{M}_{Rob}$ to some extent, as shown in Fig. 3.
@@
Figure 3: The Mask Evolution in Sec. 4.1. $\mathcal{M}_{Q}$ is obtained by filtering $\mathcal{M}_{Rob}$ from the references non-distractor regions, which is then filled by decoupling $\mathcal{M}_{D}$ and using segmentation results to get final $\mathcal{M}$ as Eq. 4 6 7. Without references filter, target regions are often misidentified as distractor.